# Patrick Kahardipraja1, Thomas Wiegand1,2,3, Wojciech Samek1,2,3, Sebastian Lapuschkin1

### v1

@@ Abstract
Recent advances in mechanistic interpretability have highlighted the potential of automating interpretability pipelines in analyzing the latent representations within LLMs. While they may enhance our understanding of internal mechanisms, the field lacks standardized evaluation methods for assessing the validity of discovered features. We attempt to bridge this gap by introducing [Uncaptioned image]FADE: Feature Alignment to Description Evaluation, a scalable model-agnostic framework for evaluating feature-description alignment. FADE evaluates alignment across four key metrics — Clarity, Responsiveness, Purity, and Faithfulness — and systematically quantifies the causes for the misalignment of feature and their description. We apply [Uncaptioned image]FADE to analyze existing open-source feature descriptions, and assess key components of automated interpretability pipelines, aiming to enhance the quality of descriptions. Our findings highlight fundamental challenges in generating feature descriptions, particularly for SAEs as compared to MLP neurons, providing insights into the limitations and future directions of automated interpretability. We release [Uncaptioned image]FADE as an open-source package at: https://github.com/brunibrun/FADE.
@@
Bruno Puri1,2,*, Aakriti Jain1,*, Elena Golimblevskaia1,*, Patrick Kahardipraja1, Thomas Wiegand1,2,3, Wojciech Samek1,2,3, Sebastian Lapuschkin1 1Department of Artificial Intelligence, Fraunhofer Heinrich Hertz Institute, Berlin, Germany 2Department of Electrical Engineering and Computer Science, Technische Universität Berlin, Berlin, Germany 3BIFOLD - Berlin Institute for the Foundations of Learning and Data, Berlin, Germany *contributed equally corresponding authors: {wojciech.samek,sebastian.lapuschkin}@hhi.fraunhofer.de
@@ 1 Introduction
Understanding the latent features of machine learning models and aligning their descriptions with human-comprehensible concepts remains a crucial challenge in AI interpretability research. Recent advances have made significant strides in this direction, by introducing automated interpretability methods Bills et al. (2023); Bykov et al. (2024); Choi et al. (2024), that leverage larger language-capable models to describe the latent representations of smaller models Bykov et al. (2023); Templeton et al. (2024); Dreyer et al. (2025). This facilitates inspection of ML models, enabling a deeper understanding of models’ behaviour. Consequently, this enhances our ability to identify or mitigate harmful responses and biases, thus improving model transparency and interpretability Lee et al. (2024); Gandikota et al. (2024).
@@
A key insight from these investigations is the highly polysemantic nature of individual neurons — they rarely correspond to single, clear concepts. This discovery has led to the development and adoption of sparse autoencoders (SAEs) McGrath et al. (2024); Bricken et al. (2023); Rajamanoharan et al. (2024), which are intended to decompose polysemantic representations by separating neuron activations into more interpretable components. While SAEs offer a promising approach for feature decomposition, their reliability remains an open question. Recent research reveals significant variability in the way SAEs capture the underlying learned features Heap et al. (2025); Paulo and Belrose (2025), thus highlighting the need for a holistic framework for the evaluation of feature-description alignment. To the best of our knowledge, there is an absence of widely accepted quantitative metrics for evaluating the quality and effectiveness of open-vocabulary feature descriptions. Different methodologies rely on custom evaluation criteria which makes it challenging to conduct meaningful, generalizable comparisons across techniques. Additionally, existing evaluation approaches typically optimize for a single metric Bills et al. (2023); Choi et al. (2024) which may not capture the full complexity of a feature’s behavior and leaves open questions about whether the model truly encodes the hypothesized concept rather than simply correlating with the measured feature. With our work, we contribute as follows:
[1] We present a robust automated evaluation framework designed for broad applicability across model architectures and their SAE implementations.[Uncaptioned image]FADE combines four metrics that allow quantitative analysis of different aspects of alignment between features and their generated descriptions. [^1] We release [Uncaptioned image]FADE as an open-source Python package (available at https://github.com/brunibrun/FADE) that includes notebooks demonstrating example neurons featured in this work.
[2] Through systematic empirical analysis, we provide insights into how various components of the autointerpretability pipeline — such as the number of layers, sample sizes, and architectural choices — affect the quality of feature descriptions.
[3] We release a selected subset of feature descriptions, presented as a part of this work for Gemma-2-2b and Gemma Scope SAEs along with their evaluations.
@@ 2 Related Work
Activation-centric methods focus on measuring how well a feature’s activations correspond to its assigned description.
One prominent approach is a simulation-based scoring, where an LLM predicts feature activations based on the description and input data, and the correlation between predicted and real activations of a feature is measured Bills et al. (2023); Bricken et al. (2023); Choi et al. (2024). While elegant, this approach can be computationally expensive and tends to favor broad, high-level explanations.
A related and conceptually more straightforward way to measure how well the description explains a feature’s behavior is to try to directly generate synthetic samples using the description and compare the resulting activations between concept and non-concept samples Huang et al. (2023); Kopf et al. (2024); Gur-Arieh et al. (2025); Shaham et al. (2025). However, generated datasets are typically small (on the order of 5–20 samples Huang et al. (2023); Gur-Arieh et al. (2025)) and often constrained to rigid syntactic structures or focus only on the occurence of particular tokens, making them less effective for evaluating abstract or open-ended language concepts Huang et al. (2023); Foote et al. (2023).
Another strategy is rating individual samples from a natural dataset for how strongly they express a concept and compare those ratings to the feature’s activations Huang et al. (2023); Paulo et al. (2024); Templeton et al. (2024).
A common limitation of activation-centric methods is that they primarily evaluate positively correlated activations while ignoring negatively encoded neurons, effectively ignoring negatively encoded features Huang et al. (2023); Kopf et al. (2024).
@@
There is a growing need for frameworks that integrate multiple perspectives to provide a comprehensive assessment of feature-to-description alignment across different interpretability methods.
For instance, prior work Bills et al. (2023); Menon et al. (2025); Gur-Arieh et al. (2025) has shown that while activation-centric and output-centric measures often correlate, they do not necessarily imply a causal relationship.
Some studies focus exclusively on SAEs McGrath et al. (2024); Paulo et al. (2024), while others analyze MLP neurons Bills et al. (2023); Choi et al. (2024). Developing an architecture-agnostic framework for feature-to-description evaluation is essential for enabling robust quantitative comparisons across interpretability approaches.
Although efforts have been made to integrate multiple evaluation perspectives Paulo et al. (2024); Gur-Arieh et al. (2025), these remain fragmented and are often too narrowly scoped to handle open-ended language descriptions.
@@ 3 Evaluating Feature Explanations
Figure 2: Refer to captionFADE can highlight different problems that arise with description generation. (a) Feature 128 contains concept of the expression “under someone’s belt”. However, the derived concept “belt”, is not clear/ specific enough to be useful for generating synthetic data, that would activate the feature. However, the description is still very close to the concept, therefore Responsiveness and Purity are high. (b) Feature 1776 strongly reacts to the word “each”, but also to many other general, unrelated words, resulting in lower Clarity and Purity. (c) The description for feature 7657 catches its main concept, low Purity indicates that this feature is clearly polysemantic. (d) Description for feature 10647 is expressed too broad, resulting in low Responsiveness and Purity.
@@ 4 Experiments
In this section, we apply [Uncaptioned image]FADE to assess the quality of descriptions generated for various state-of-the-art feature descriptions Choi et al. (2024); Lieberum et al. (2024). Our goal is to demonstrate that the proposed framework provides a robust, multidimensional measure of feature to feature description alignment.
@@
As a natural dataset for the evaluations we use samples drawn from the test partition of the Pile dataset Gao et al. (2020), preprocessed as shown in Appendix B. As evaluating LLM we use the OpenAI model gpt-4o-mini-2024-07-18 unless stated otherwise. Prompts for the evaluating LLM as well as details on the hyperparameters can be found in Appendix D.2. We run the experiments on $10^{3}$ randomly chosen features from a single layer of a model: layer 20 for Gemma-2-2b Riviere et al. (2024), layer 20 of Gemma Scope SAEs Lieberum et al. (2024), and layer 19 of Llama-3.1-8B-Instruct Dubey et al. (2024) (see Appendix C.1 for details). The evaluation results have a high variance, which is caused by both the inherent difficulty of interpreting some features as well as the quality of the ascribed feature descriptions. Therefore the mean values are demonstrated only if they aid the analysis of metrics distributions. For all of the presented tables we demonstrate the full distributions as kernel-density estimations with bandwidth adjustment factor 0.3 in Appendix 6.
@@
Feature descriptions, which we refer to as MaxAct*, are generated based on samples of the train partition of the Pile dataset, that demonstrate maximum activation on the feature, similarly to methods utilized in Bills et al. (2023); Paulo et al. (2024); Rajamanoharan et al. (2024), that we refer to as MaxAct. The minor differences between MaxAct and MaxAct* are prompts, optimized on qualitative analysis provided via [Uncaptioned image]FADE, and preprocessing steps of the dataset. The automated interpretability pipeline is described in Appendix C.1.
@@ 4.1 Depth and Reliability of Evaluations
As the evaluating LLM is one of the most computationally expensive components of our framework, selecting a model that balances performance and cost is critical. Larger models generally achieve better performance, but at significantly higher computational costs. To determine a minimal feasible model size and capability required for effective evaluation, we conduct a quantitative analysis of concept-expression ratings across various open-weight and proprietary models, using GPT-4o as a baseline due to its superior benchmark performance OpenAI (2024). We evaluate models on Neuronpedia Lin (2023) feature descriptions for the Gemma Scope SAEs, generated via MaxAct method Rajamanoharan et al. (2024). By comparing deviations in concept strength ratings between GPT-4o and other models, we assess their relative performance (see Table 1).
@@
Table 1: Concept rating procedure for different evaluating LLMs. The GPT-4o baseline shows the number of occurrences per class. The other models show their alignment with the GPT-4o rating in percent. The “Valid” column shows the percentage of samples that were correctly classified. Class 0 represents no alignment with the concept, class 1 a partial alignment and class 2 means the samples clearly exhibits the concept.
@@
To compare MLP neurons and SAE features, we analyze Gemma-2-2b and Gemma Scope SAEs. While Gemma-2 outperforms Gemma Scope SAEs in average clarity (see Table 4), the clarity score distribution reveals a left-skewed peak for SAEs, as depicted on Figure 4. Further analysis identifies a cluster of features with low clarity but moderate to high responsiveness and purity. These features have descriptions that approximate the encoded concept but lack the precision to strongly activate the SAE feature (see (d) in Figure 2). This suggests that despite greater monosemanticity, interpreting SAE features remains challenging due to the difficulty of generating precise descriptions. In contrast, responsiveness and purity are higher on average for SAEs, as these metrics are less sensitive to imprecise descriptions and still align with the underlying concept. The higher purity in SAEs aligns with their increased monosemanticity.
@@ 4.2 Evaluating Autointerpretability: Prompts, Examples, and Model Size
Similarly to the experiment, comparing different evaluation models presented in Table 1, we compare different explainer models and demonstrate the results in Table 7. GPT-4o achieves the highest scores, with Llama-3.3-70B-Instruct (AWQ 4-bit quantized) and GPT-4o mini as close alternatives. The smaller models struggle with assigning reasonable feature descriptions, in particularly Llama-3.2-1B, which frequently fails to maintain a consistent response structure (see Appendix 6).
@@ 5 Conclusion
Despite presenting a comprehensive and robust evaluation framework, our work has certain limitations that we want to highlight here: One key limitation is the potential biases in the LLMs used for both rating and synthetic data generation. These biases can affect the evaluation process and perpetuate biases, especially in automated interpretability. For instance, an LLM might recognize a feature encoding a concept in English as directly representing that concept, whereas the same feature in another language might be classified with the additional specification of the language. This discrepancy could lead to unintended biases when steering models based on these interpretations. Similar issues may arise from biases present in the pre-training datasets used in our evaluation procedure. Another limitation is related to the steering behavior in the faithfulness pipeline. Our current implementation does not explicitly verify whether the generated sequences under modification remain grammatically correct and semantically meaningful. However, minor modifications to the prompt could potentially address this issue in the future. Finally, our faithfulness measure is not well-suited for handling inhibitory neurons. A neuron may causally inhibit the presence of a concept in a model’s output, but our metric, by design, does not effectively capture decreases in the appearance of sparse concepts. This limitation arises both from the definition of our faithfulness metric and the inherent challenges in measuring such suppression effects for already sparse concepts.

### v2

# Thomas Wiegand1,2,3, Wojciech Samek1,2,3, Sebastian Lapuschkin1,4

@@ Abstract
Recent advances in mechanistic interpretability have highlighted the potential of automating interpretability pipelines in analyzing the latent representations within LLMs. While this may enhance our understanding of internal mechanisms, the field lacks standardized evaluation methods for assessing the validity of discovered features. We attempt to bridge this gap by introducing [Uncaptioned image]FADE: Feature Alignment to Description Evaluation, a scalable model-agnostic framework for automatically evaluating feature-to-description alignment. FADE evaluates alignment across four key metrics – Clarity, Responsiveness, Purity, and Faithfulness – and systematically quantifies the causes of the misalignment between features and their descriptions. We apply [Uncaptioned image]FADE to analyze existing open-source feature descriptions and assess key components of automated interpretability pipelines, aiming to enhance the quality of descriptions. Our findings highlight fundamental challenges in generating feature descriptions, particularly for SAEs compared to MLP neurons, providing insights into the limitations and future directions of automated interpretability. We release [Uncaptioned image]FADE as an open-source package at: https://github.com/brunibrun/FADE.
@@
Bruno Puri1,2,*, Aakriti Jain1,*, Elena Golimblevskaia1,*, Patrick Kahardipraja1, Thomas Wiegand1,2,3, Wojciech Samek1,2,3, Sebastian Lapuschkin1,4 1Department of Artificial Intelligence, Fraunhofer Heinrich Hertz Institute 2Department of Electrical Engineering and Computer Science, Technische Universität Berlin 3BIFOLD - Berlin Institute for the Foundations of Learning and Data 4Centre of eXplainable Artificial Intelligence, Technological University Dublin corresponding authors: {wojciech.samek,sebastian.lapuschkin}@hhi.fraunhofer.de
**footnotetext: These authors contributed equally.
@@ 1 Introduction
Understanding the latent features of machine learning models and aligning their descriptions with human-comprehensible concepts remains a crucial challenge in AI interpretability research. Recent advances have made significant strides in this direction, by introducing automated interpretability methods Bills et al. (2023); Bykov et al. (2024); Choi et al. (2024) that leverage larger language-capable models to describe the latent representations of smaller models Bykov et al. (2023); Templeton et al. (2024); Dreyer et al. (2025). This facilitates inspection of ML models, enabling a deeper understanding of models’ behavior which enhances our ability to identify or mitigate harmful responses and biases Lee et al. (2024); Gandikota et al. (2024).
@@
A key insight from these investigations is the highly polysemantic nature of individual neurons – they rarely correspond to single, clear concepts. This discovery has led to the development and adoption of sparse autoencoders (SAEs) McGrath et al. (2024); Bricken et al. (2023); Rajamanoharan et al. (2024), which are intended to decompose polysemantic representations by separating neuron activations into more interpretable components. While SAEs offer a promising approach for feature decomposition, their reliability remains an open question. Recent research reveals significant variability in the way SAEs capture the underlying learned features Heap et al. (2025); Paulo and Belrose (2025), thus highlighting the need for a holistic framework for the evaluation of feature-to-description alignment. To the best of our knowledge, there is an absence of widely accepted quantitative metrics for evaluating the quality and effectiveness of open-vocabulary feature descriptions. Different methodologies rely on custom evaluation criteria which makes it challenging to conduct meaningful, generalizable comparisons across techniques. Additionally, existing evaluation approaches typically optimize for a single metric Bills et al. (2023); Choi et al. (2024) which may not capture the full complexity of a feature’s behavior and leaves open questions about whether the model truly encodes the hypothesized concept rather than simply correlating with the measured feature. With our work, we contribute as follows:
[1] We present a robust automated evaluation framework designed for broad applicability across model architectures and their SAE implementations.[Uncaptioned image]FADE combines four metrics that allow quantitative analysis of different aspects of alignment between features and their generated descriptions. [^1] We release [Uncaptioned image]FADE as an open-source Python package available at https://github.com/brunibrun/FADE, and include example notebooks as well as some of the feature evaluations presented in this work.
[2] Through systematic empirical analysis, we provide insights into how various components of the autointerpretability pipeline – such as the number of layers, sample sizes, and architectural choices – affect the quality of feature descriptions.
[3] We release a set of feature descriptions, presented as part of this work for Gemma-2-2b and Gemma Scope SAEs along with their evaluations.
@@ 2 Related Work
Activation-centric methods focus on measuring how well a feature’s activations correspond to its assigned description. One prominent approach is a simulation-based scoring, where an LLM predicts feature activations based on the description and input data, and the correlation between predicted and real activations of a feature is measured Bills et al. (2023); Bricken et al. (2023); Choi et al. (2024). While elegant, this approach can be computationally expensive and tends to favor broad, high-level explanations. A related and conceptually more straightforward way to measure how well the description explains a feature’s behavior is to try to directly generate synthetic samples using the description and compare the resulting activations between concept and non-concept samples Huang et al. (2023); Kopf et al. (2024); Gur-Arieh et al. (2025); Shaham et al. (2025). However, generated datasets are typically small (on the order of 5–20 samples Huang et al. (2023); Gur-Arieh et al. (2025)) and often constrained to rigid syntactic structures or focus only on the occurrence of particular tokens, making them less effective for evaluating abstract or open-ended language concepts Huang et al. (2023); Foote et al. (2023). Another strategy is rating individual samples from a natural dataset for how strongly they express a concept and compare those ratings to the feature’s activations Huang et al. (2023); Paulo et al. (2024); Templeton et al. (2024). A common limitation of activation-centric methods is that they primarily assume activations to be positively correlated with the concept, thus ignoring negatively encoded features Huang et al. (2023); Kopf et al. (2024).
@@
There is a growing need for frameworks that integrate multiple perspectives to provide a comprehensive assessment of feature-to-description alignment across different interpretability methods. For instance, prior work Bills et al. (2023); Menon et al. (2025); Gur-Arieh et al. (2025) has shown that while activation-centric and output-centric measures often correlate, they do not necessarily imply a causal relationship. Some studies focus exclusively on SAEs McGrath et al. (2024); Paulo et al. (2024), while others analyze MLP neurons Bills et al. (2023); Choi et al. (2024). Developing an architecture-agnostic framework for feature-description evaluation is essential for enabling robust quantitative comparisons across interpretability approaches. Although efforts have been made to integrate multiple evaluation perspectives Paulo et al. (2024); Gur-Arieh et al. (2025), these remain fragmented and are often too narrowly scoped to handle open-ended language descriptions.
@@ 3 Evaluating Feature Explanations
Figure 2: Refer to captionFADE can highlight different problems that arise with description generation. (a) The description for feature 7657 catches its main concept, low Purity indicates that this feature is clearly polysemantic. (b) Feature 1776 strongly reacts to the word “each”, but also to many other general, unrelated words, resulting in lower Clarity and Purity. (c) Description for feature 10647 is expressed too broad, resulting in low Responsiveness and Purity. (d) Feature 128 encodes the concept “under someone’s belt”. However, the derived description “belt”, is not clear or specific enough to be useful for generating synthetic data that would activate the feature, but it is still closely related to the concept, therefore Responsiveness and Purity are high.
@@ 3.2 Evaluation Metrics
3.3 [Uncaptioned image]FADE Evaluation Framework
FADE is designed to work with a wide variety of Subject Models, including most HuggingFace Transformers (Wolf et al., 2020) as well as any feature implemented as a named module in PyTorch. Moreover, it is extensible: interpretability tools such as SAEs or various supervised interpretability techniques can be integrated with minimal effort, provided they implement some basic steering functionality. In addition, we offer an interface for a diverse set of evaluating LLMs, whether open-weight or proprietary, with support for frameworks such as vLLM, Ollama, OpenAI, and Azure APIs.
@@ 4 Experiments
In this section, we apply [Uncaptioned image]FADE to assess the alignment of features and their descriptions generated via various state-of-the-art automated interpretability methods Choi et al. (2024); Lieberum et al. (2024). Our goal is to demonstrate that the proposed framework provides a robust, multidimensional measure of feature-to-description alignment.
@@
As a natural dataset for the evaluations we use samples drawn from the test partition of the Pile dataset Gao et al. (2020), preprocessed as shown in Appendix B, giving us approximately 5 million samples. As evaluating LLM we use the OpenAI model gpt-4o-mini-2024-07-18 unless stated otherwise. Prompts for the evaluating LLM as well as details on the hyperparameters can be found in Appendix D.2. We run the experiments on $10^{3}$ randomly chosen features from a single layer of a model: layer 20 for Gemma-2-2b Riviere et al. (2024), layer 20 of Gemma Scope SAEs Lieberum et al. (2024), and layer 19 of Llama-3.1-8B-Instruct Dubey et al. (2024) (see Appendix C.1 for details). The evaluation results have a high variance, which is caused by both the inherent difficulty of interpreting some features as well as the quality of the ascribed feature descriptions. Therefore the mean values are reported only where they help interpret the distribution of the metrics. For all of the presented tables we demonstrate the full distributions as kernel-density estimations with bandwidth adjustment factor 0.3 in Appendix 7.
@@
Feature descriptions, which we refer to as MaxAct*, are generated based on samples of the train partition of the Pile dataset, that demonstrate maximum activation on the feature, similarly to methods utilized in Bills et al. (2023); Paulo et al. (2024); Rajamanoharan et al. (2024). The minor differences between MaxAct and MaxAct* are prompts, optimized on qualitative analysis provided via [Uncaptioned image]FADE, and preprocessing steps of the dataset. The automated interpretability pipeline is described in Appendix C.1.
@@ 4.1 Depth and Reliability of Evaluations
As the evaluating LLM is one of the most computationally expensive components of our framework, selecting a model that balances performance and cost is critical. Larger models generally achieve better performance, but at significantly higher computational costs. To determine a minimal feasible model size and capability required for effective evaluation, we conduct a quantitative analysis of concept-expression ratings across various open-weight and proprietary models, using GPT-4o as a baseline due to its superior benchmark performance OpenAI (2024). We evaluate models using Neuronpedia Lin (2023) feature descriptions for the Gemma Scope SAEs, generated via the MaxAct method Rajamanoharan et al. (2024). By comparing deviations in concept strength ratings between GPT-4o and other models, we assess their relative performance (see Table 1).
@@
Table 1: Concept rating procedure for different evaluating LLMs. The GPT-4o baseline shows the number of occurrences per class. The other models show their alignment with the GPT-4o rating in percent. The “Valid” column shows the percentage of samples that were correctly classified. Class 0 represents no alignment with the concept, class 1 a partial alignment and class 2 means the samples clearly exhibit the concept.
@@
Figure 5: Examples of feature descriptions, obtained via MaxAct*, demonstrating low Clarity, but medium or high Responsiveness and Purity. Tokens are highlighted in green when the feature is activated, and in yellow if the feature is not activated even though the concept is present. Feature 475: “Frequent presence of token ’self’ indicating object oriented programming concept”. Feature 653: “Word ’make’ in different forms, expressions and concepts”. Feature 821: “Expressions indicating lists or explanations”.
@@
To compare MLP neurons and SAE features, we analyze Gemma-2-2b and Gemma Scope SAEs. While Gemma-2 outperforms Gemma Scope SAEs in average Clarity (see Table 4), the Clarity score distribution reveals a left-skewed peak for SAEs, as depicted in Figure 4. Further analysis identifies a cluster of features with low Clarity but moderate to high Responsiveness and Purity. These features have descriptions that approximate the encoded concept but lack the precision to strongly activate the SAE feature. Figure 5 presents the heatmap for feature 475, whose description emphasizes the occurrence of the token “self.” However, the feature does not activate on all instances of “self” (highlighted in yellow), indicating that a crucial aspect of the concept is missing or remains unclear. Additionally, the feature activates on other tokens, further suggesting that the description is incomplete. This is reflected in the evaluation metrics: Clarity of 0.005 indicates that the concept is not expressed precisely enough for the evaluating LLM to generate synthetic data that reliably activates the feature. A Responsiveness score of 0.93 suggests that the feature does activate on natural data aligned with the concept, while a Purity score of 0.81 reveals that although the feature is primarily associated with the described concept, it also responds to other inputs. Similar issues arise in features 653 and 821, where the underlying concepts appear highly specific – activating on a particular token within a specific context. However, their descriptions are overly broad, making it difficult to generate synthetic data that reliably triggers the feature.
This suggests that despite greater monosemanticity, interpreting SAE features remains challenging due to the difficulty in generating precise descriptions. In contrast to the Clarity metric, Responsiveness and Purity are on average higher for SAEs, as these metrics are less sensitive to imprecise descriptions and still align with the underlying concept. The higher Purity in SAEs aligns with our expectations of their greater monosemanticity.
@@ 4.2 Evaluating Autointerpretability: Prompts, Examples, and Model Size
Similarly to the experiment, comparing different Evaluation Models presented in Table 1, we compare different Explainer Models and demonstrate the results in Table 7. GPT-4o achieves the highest scores, with Llama-3.3-70B-Instruct (AWQ 4-bit quantized) and GPT-4o mini as close alternatives. The smaller models struggle with assigning reasonable feature descriptions, in particular Llama-3.2-1B, which frequently fails to maintain a consistent response structure (see Appendix 7).
@@ 5 Conclusion
Despite presenting a comprehensive and robust evaluation framework, our work has certain limitations that we want to highlight here: One key limitation when using LLMs is the inherent biases that can influence both rating and synthetic data generation, thus affecting our evaluation metrics. When rating concepts, an LLM’s inherent bias may systematically lead to lower ratings for certain cultural contexts, domains, or languages outside its primary training distribution. Similarly, during synthetic data generation, LLMs may produce less diverse or representative data for underrepresented domains and low-resource languages. For example, an LLM might recognize a feature encoding a concept in English as directly representing that concept, whereas the same feature in another language might be classified with the additional specification of the language. This discrepancy could also lead to unintended biases when steering models based on these interpretations.
Similar issues may arise from biases present in the datasets used in the evaluation procedure. While we used The Pile dataset due to its extensive nature, underrepresented concepts may still lead to less stable purity and responsiveness measures. Although FADE can help flag such cases and enable us to achieve fuller coverage by identifying the aforementioned training dataset issues, ultimately a more balanced corpus would be beneficial.
Another limitation is related to the steering behavior in the Faithfulness pipeline. Our current implementation does not explicitly verify whether the generated sequences under modification remain grammatically correct and semantically meaningful. However, this can be mitigated in the future by integrating KL-divergence to control the magnification factor, used for steering the features Paulo et al. (2024); Gur-Arieh et al. (2025). Finally, our Faithfulness measure is not well-suited for capturing the behavior of inhibitory neurons. While a neuron may causally suppress the presence of a concept in the model’s output, our metric, by design, is limited in detecting decreases in the activation of sparse concepts. Moreover, current automated interpretability methods, which our evaluation framework is intended to support, typically do not focus on identifying inhibitory neurons. However, if an inhibitory feature is redefined in affirmative terms (e.g., instead of "the feature suppresses English," phrased as "the feature promotes non-English languages"), our Faithfulness metric can handle such cases appropriately. Nonetheless, a more comprehensive investigation into inhibitory neurons, both in terms of automated interpretability and evaluation, remains necessary.